40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 4 20:03:38 2012, Jenne, Update, General, in-vac progress 
    Reply  Tue Sep 4 20:20:47 2012, jamie, Update, General, problematic tip-tilts 
Message ID: 7338     Entry time: Tue Sep 4 20:03:38 2012     Reply to this: 7341
Author: Jenne 
Type: Update 
Category: General 
Subject: in-vac progress 

Opened ETMY, beam was high and to the right (if you look at the face of ETMY).  Tried walking beam up, since just doing PZT2 caused clipping at the BS before we got to the correct spot on ETMY.  Moved PZT1, then PZT2, to translate beam, but we couldn't get far enough without starting to fall off of PZT2.  Put PZT1 approx. back where it was. Jamie tapped on the top of PR3 (tip tilt just before BS), and then did some compensation with PZT2 to get the beam through the BS target to the center of ETMY.

The beam is very dim at ETMY.  We ended up holding the big IR card with holes such that one of the holes was near the center of the optic, in front of the cage.  Then one person turned off all the room lights so we could see the beam, another person moved PZT2 and PR3, then lights person turned on and off the lights so we could compare beam position with hole position.  A total pain, but it ended up working better than just trying to follow the beam with a card.

We clearly need a better plan for adjusting the tip tilts in pitch, because utilizing their hysteresis is ridiculous.  Koji and Steve are thinking up a set of options, but so far it seems as though all of those options should wait for our next "big" vent.  So for now, we have just done alignment by poking the tip tilt.

Tomorrow, we want to open up the MC doors, open up ETMY, and look to see where the beam is on the optic.  I am concerned that the hysteresis will relax over a long ( >1hour ) time scale, and we'll loose our pointing.  After that, we should touch the table enough to trip the BS, PRM optics, since Koji is concerned that perhaps the tip tilt will move in an earthquake.  Jamie mentioned that he had to poke the tip tilt a pretty reasonable amount to get it to change a noticeable amount at ETMY, so we suspect that an earthquake won't be a problem, but we will check anyway.

After doing all of that, we adjusted IPANG so that the beam gets out of the vacuum envelope.  The beam must squeeze between the wall of the black beam tube and a lens for the oplev, so there is a very, very limited amount of space.  The eventual new tables will be very helpful for alleviating this, but for now we must live with it.  Even though it is pretty squishy right at the edge of the table, with the new layout we should think about giving IPANG a little more space.  Basically, if there is any Yaw motion, the beam going to the QPD will be clipped, and we may get confusing info.  We moved some of the IPANG optics that are on the end transmission table so the beam is centered on the optics while it just makes it through the space between the wall and the lens.  The spot was centered on the IPANG QPD. 

We still need to check on IPPOS, but it is always easier than IPANG.

General notes:

We discovered that the 45deg SOS beam targets are awesome, especially the ones with the irises.  The plain hole ones have very small holes relative to our beam size, so they are much more useful for the MC optics (which they were designed for).  The 45deg targets were made such that the target holes do not line up with the mounting holes.  This is good, since the mounting holes are lower than the center of the optic.  I don't think ericQ and I realized that on Friday, so it's probable that we had installed the target upside-down.  We still need to remake the 0deg targets for the next vent.

Jamie dogged down the new 'bathroom mirror' that lets us see BS and PRM on the same camera view.  He also adjusted some of the masses on the BS table to relevel the table. We need to (at some point) rename the AS_SPARE camera to something like BS_PRM, since we plugged the new camera into the AS_SPARE port on the videoswitch.

tl;dr: Input beam adjusted so we're hitting center of ETMY.  IPANG coming out of vacuum, QPD centered. Need better tip tilt in-situ adjustment capability.

ELOG V3.1.3-